OpenStudy (anonymous):

If y = –5 when x = 25, find x when y = 45

OpenStudy (lgbasallote):

@lisa-sws i think you forgot to mention if this is directly or inversely proportional..

OpenStudy (anonymous):

direct @Igbasallote

OpenStudy (anonymous):

the relationship is easy: \[x= y ^{2}\] so if y=45 \[x=45^{2} = 2025\]

OpenStudy (lgbasallote):

it's directly proportional y = kx when y = -5 then x = 25 -5 = 25k -5/25 = lk -1/5 = k what is x when y = 45 y = kx 45 = -1/5 x multiply both sides by -5 45*-5 = x simplify that to get your final answer

OpenStudy (lgbasallote):

i would suggest you post the FULL question to be clear @lisa-sws

OpenStudy (anonymous):

Write a direct variation equation that relates x and y. Assume that y varies directly as x. Then solve. If y = –5 when x = 25, find x when y = 45 @Igbasallote

OpenStudy (lgbasallote):

aahh then use the one i gave you

OpenStudy (anonymous):

thanks.